OpenStudy (anonymous):

The perimeter of a square with side x-1 units is 3x+2 units. Find its numerical value of its area. a. 20 b. 24 c. 25 d. 36

OpenStudy (anonymous):

is that -1

hartnn (hartnn):

perimeter of square = 4* side! so, 3x+2 = 4(x-1) can you solve this for x ?

hartnn (hartnn):

then AREA is just side^2 = (x-1)^2 but first get the value of x

OpenStudy (anonymous):

Is x=6? Then, what's next? Get the area?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first substitutes the value of x

hartnn (hartnn):

yes, x=6, so x-1 = 5 area = 5^2 =... ?

OpenStudy (anonymous):

25!! Yehey! Thanks! :)

hartnn (hartnn):

correct! welcome ^_^